Circular No. 7192

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

V382 VELORUM S. N. Shore, Indiana University, South Bend; H. E. Bond and R.

Downes, Space Telescope Science Institute; S. Starrfield, Arizona State University; R. D. Gehrz, University of Minnesota; J. Krautter, Landessternwarte Heidelberg; and C. E. Woodward, University of Wyoming, write: “We observed V382 Vel with the Hubble Space Telescope’s Imaging Spectrograph (STIS) on May 31.3 UT, using the medium-resolution échelle gratings (E140M and E230M). Three exposures were taken, centered at 142.5, 197.8, and 270.7 nm with resolutions of ∼ 45 000 (E140M) and 30 000 (E230M). Mg II h and k emission was detected with FWZI = 4000 km/s, a maximum blueshift of −2000 km/s, and an integrated line flux of 8.2 × 10^{−9} erg cm^{−2} s^{−1} (no reddening correction); the line profile shows neither P-Cyg absorption nor fine structure but is steeper on the blueward emission side. The only other obvious features are the iron-peak absorption bands that dominate the short-wavelength (< 200 nm) spectrum, and the Al II and Fe-peak blend in emission at 265 nm. The width of the Mg II line, which is nearly optically thin, and the resemblence to the spectrum of V1974 Cyg at a similar stage of the outburst, support the suggestion (IAUC 7185) that this is an ONeMg-type nova. The interstellar Mg II lines show a broad, completely optically thick component with mean radial velocity +24.2 km/s and an optically thin narrower component at +48.5 km/s (the Mg II 279.6-nm/280.3-nm line ratio for the latter component is ∼ 2.4). The integrated flux from 120.0 to 320.0 nm is 9.6 × 10^{−8} erg cm^{−1} s^{−1}, or 3.7 × 10^{−7} if corrected for reddening of E(B − V) = 0.2. The reddening is probably low, as there is no evidence for the 217.5-nm interstellar feature. An additional STIS observation will be made on June 21.6. Continued spectroscopy at all wavelengths is urged.”

COMET C/1999 H1 (LEE) Total visual magnitude estimates: May 17.07 UT, 6.5 (J. E. Bortle,

Stormville, NY, 10×50 binoculars); 31.33, 6.6 (D. A. J. Seargent, The Entrance, N.S.W., 10×50 binoculars); June 2.31, 6.9 (S. T. Rae, Palmerston North, New Zealand, 10×50 binoculars); 4.85, 6.4 (R. Haver, Frasso Sabino, Rieti, Italy, 15×80 binoculars); 7.44, 6.4 (A. Pearce, Nedlands, W. Australia, 20×80 binoculars).

1999 June 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green